LINE POSITIONS AND INTENSITIES OF PHOSPHINE IN THE 3 μm REGION FOR PLANETARY AP-PLICATIONS

\underline{L. SAGUI}, I. KLEINER, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay Cedex, France; R. A. H. BUTLER and L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA.

Seven absorption bands of PH_{3} at 3 μm, i.e 3ν_{2}, 2ν_{2} + ν_{4}, 2ν_{4} + ν_{2}, ν_{1} + ν_{2}, ν_{3} + ν_{2}, ν_{3} + ν_{4} and ν_{1} + ν_{4} were simultaneously analysed from spectra recorded at a resolution of 0.012 cm^{−1} using the Fourier transform spectrometer located at the Kitt Peak Observatory^{a}. A theoretical model suited to the strong vibro-rotational couplings among the seven bands was used. Using ground state combination differences, several hundred transitions have been assigned for these seven bands in the range 2750–3600 cm^{−1}. At present a set of 33 energy parameters was required to reproduce the experimental wavenumbers for six bands with an overall standard deviation of 0.1 cm^{−1} up to J = 9. 1355 lines were also measured and their analysis is underway. The final objective is to provide a complete prediction of line positions and intensities of Phosphine at 3 μm for remote sensing of Saturn.

aPart of the research reported in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.